# Identification of Dynamic Nonlinear Panel Models

### v2

@@ A.3 Proof of Proposition 2 and 3
$1-\Pr(Y_{i1}=0,z_{1}^{\prime}\beta_{0}+X_{i1}^{\prime}\gamma_{0}\geq a\mid z_{12})\geq\Pr(Y_{i2}=1,z_{2}^{\prime}\beta_{0}+X_{i2}^{\prime}\gamma_{0}\leq a\mid z_{12}),$
for $a\in\{z_{1}^{\prime}\beta_{0}+a_{1}^{\prime}\gamma_{0},...,z_{1}^{\prime}\beta_{0}+a_{K}^{\prime}\gamma_{0},z_{2}^{\prime}\beta_{0}+a_{1}^{\prime}\gamma_{0},...,z_{2}^{\prime}\beta_{0}+a_{K}^{\prime}\gamma_{0}\}$, and another identifying condition switches the order of period 1 and 2.
Let $a=z_{1}^{\prime}\beta_{0}+a_{k}^{\prime}\gamma_{0}$, [^11] The value of $a=z_{2}^{\prime}\beta_{0}+a_{k}^{\prime}\gamma_{0}$ leads to the same identifying condition., then the above upper bound can be further bounded as
$1-\Pr(Y_{i1}=0,z_{1}^{\prime}\beta_{0}+X_{i1}^{\prime}\gamma_{0}\geq z_{1}^{\prime}\beta_{0}+a_{k}^{\prime}\gamma_{0}\mid z_{12})\leq 1-\Pr(Y_{i1}=0,X_{i1}=a_{k}\mid z_{12}).$
When $z_{1}^{\prime}\beta_{0}-z_{2}^{\prime}\beta_{0}\geq 0$ which implies $z_{1}^{\prime}\beta_{0}+a_{k}^{\prime}\gamma_{0}\geq z_{2}^{\prime}\beta_{0}+a_{k}^{\prime}\gamma_{0}$, then the lower bound can be bounded below as
$\Pr(Y_{i2}=1,z_{2}^{\prime}\beta_{0}+X_{i2}^{\prime}\gamma_{0}\leq z_{1}^{\prime}\beta_{0}+a_{k}^{\prime}\gamma_{\mid}z_{12})\leq\Pr(Y_{i2}=1,X_{i2}=a_{k}\mid z_{12}).$
@@
If $\text{If }z_{1}^{\prime}\beta_{0}-z_{2}^{\prime}\beta_{0}\geq 0\Longrightarrow 1-\Pr(Y_{i1}=0,X_{i1}=a_{k}\mid z_{12})\geq\Pr(Y_{i2}=1,X_{i2}=a_{k}\mid z_{12}).$
@@
$01-\Pr(Y_{i1}=0,X_{i1}=a_{k}\mid z_{12})<\Pr(Y_{i2}=1,X_{i2}=a_{k}\mid z_{12})\Longrightarrow\Delta z^{\prime}\beta_{0}>$ 0 .
@@
$01-\Pr(Y_{i1}=1,X_{i1}=a_{k}\mid z_{12})<\Pr(Y_{i2}=0,X_{i2}=a_{k}\mid z_{12})\Longrightarrow\Delta z^{\prime}\beta_{0}<$ 0 .
@@
$(z_{1},z_{2})\in\mathcal{Z}^{j}_{3}\ \Longrightarrow\ (x_{1}^{j}-x_{2}^{j})\gamma^{j}_{0}<\Delta z^{\prime}\beta_{0},$
$(z_{1},z_{2})\in\mathcal{Z}^{j}_{4}\ \Longrightarrow\ (x_{1}^{j}-x_{2}^{j})\gamma^{j}_{0}>\Delta z^{\prime}\beta_{0}.$
As previously shown, when $(z_{1},z_{2})\in\mathcal{Z}_{2}$, it implies that $\Delta z^{\prime}\beta_{0}<0$. Therefore, when $(z_{1},z_{2})\in\mathcal{Z}_{2}\cap\mathcal{Z}^{j}_{3}$, we have $(x_{1}^{j}-x_{2}^{j})\gamma^{j}_{0}<\Delta z^{\prime}\beta_{0}<0$ and the sign of $\gamma^{j}_{0}$ is identified given $x_{1}^{j}\neq x_{2}^{j}$. Similarly, when $(z_{1},z_{2})\in\mathcal{Z}_{1}\cap\mathcal{Z}^{j}_{4}$, the sign of $\gamma^{j}_{0}$ is also identified given $(x_{1}^{j}-x_{2}^{j})\gamma^{j}_{0}>\Delta z^{\prime}\beta_{0}>0$. Proposition 3 requires that for any $j\leq d_{x}$, either $\mathcal{Z}_{2}\cap\mathcal{Z}^{j}_{3}\neq\emptyset$ or $\mathcal{Z}_{1}\cap\mathcal{Z}^{j}_{4}\neq\emptyset$ so that the sign of $\gamma_{0}^{j}$ is identified for any $j$.
A.4 Relationship to Khan, Ponomareva, and Tamer (2023)
Lemma 3.
Under Assumption 1 and $X_{it}=Y_{i,t-1}$, condition (7) implies the identifying restriction Khan, Ponomareva, and Tamer (2023), presented as follows:
$(i)$ $P\left(y_{it}=1\mid z_{st}\right)>P\left(y_{is}=1\mid z_{st}\right)\Longrightarrow\left(z_{t}-z_{s}\right)^{\prime}\beta_{0}+|\gamma_{0}|>0,$
$(ii)$ $P\left(y_{it}=1\mid z_{st}\right)>1-P\left(y_{i,s-1}=1,y_{is}=0\mid z_{st}\right)\Longrightarrow\left(z_{t}-z_{s}\right)^{\prime}\beta_{0}-\min\{0,\gamma_{0}\}>0,$
$(iii)$ $P\left(y_{it}=1\mid z_{st}\right)>1-P\left(y_{i,s-1}=0,y_{is}=0\mid z_{st}\right)\Longrightarrow\left(z_{t}-z_{s}\right)^{\prime}\beta_{0}+\max\{0,\gamma_{0}\}>0,$
$(iv)$ $P\left(y_{i,t-1}=1,y_{it}=1\mid z_{st}\right)>P\left(y_{is}=1\mid z_{st}\right)\Longrightarrow\left(z_{t}-z_{s}\right)^{\prime}\beta_{0}+\max\{0,\gamma_{0}\}>0,$
$(v)$ $P\left(y_{i,t-1}=1,y_{it}=1\mid z_{st}\right)>1-P\left(y_{i,s-1}=1,y_{is}=0\mid z_{st}\right)\Longrightarrow\left(z_{t}-z_{s}\right)^{\prime}\beta_{0}>0,$
$(vi)$ $P\left(y_{i,t-1}=1,y_{it}=1\mid z_{st}\right)>1-P\left(y_{i,s-1}=0,y_{is}=0\mid z_{st}\right)\Longrightarrow\left(z_{t}-z_{s}\right)^{\prime}\beta_{0}+\gamma_{0}>0,$
$(vii)$ $P\left(y_{i,t-1}=0,y_{it}=1\mid z_{st}\right)>P\left(y_{is}=1\mid z_{st}\right)\Longrightarrow\left(z_{t}-z_{s}\right)^{\prime}\beta_{0}-\min\{0,\gamma_{0}\}>0,$
$(viii)$ $P\left(y_{i,t-1}=0,y_{it}=1\mid z_{st}\right)>1-P\left(y_{i,s-1}=1,y_{is}=0\mid z_{st}\right)\Longrightarrow\left(z_{t}-z_{s}\right)^{\prime}\beta_{0}-\gamma_{0}>0,$
$(ix)$ $P\left(y_{i,t-1}=0,y_{it}=1\mid z_{st}\right)>1-P\left(y_{i,s-1}=0,y_{is}=0\mid z_{st}\right)\Longrightarrow\left(z_{t}-z_{s}\right)^{\prime}\beta_{0}>0.$
@@ A.4 Relationship to Khan, Ponomareva, and Tamer (2023)
When the endogenous regressor is the lagged dependent variable $X_{it}=Y_{it-1}$, the identifying condition in Proposition 1 is given as
$U_{s}(a):=1-\Pr(Y_{is}=0,z_{s}^{\prime}\beta_{0}+Y_{i,s-1}\gamma_{0}\geq a\mid z_{st})\\ \geq\Pr(Y_{it}=1,z_{t}^{\prime}\beta_{0}+Y_{i,t-1}\gamma_{0}\leq a\mid z_{st}):=L_{t}(a).$
Since $Y_{i,t-1}$ only takes two values $Y_{i,t-1}\in\{0,1\}$, we can express the probability in the above bounds as a mixture of $Y_{i,t-1}=1$ and $Y_{i,t-1}=0$:
$U_{s}(a)=1-\Pr(Y_{is}=0,Y_{i,s-1}=1\mid z_{st})\mathbf{\mathbbm{1}}\{z_{s}^{\prime}\beta_{0}$ $+\gamma_{0}\geq a\}$
$-\Pr(Y_{is}=0,Y_{i,s-1}=0\mid z_{st})\mathbf{\mathbbm{1}}\{z_{s}^{\prime}\beta_{0}\geq a\}.$
Similarly, the lower bound is given as
$L_{t}(a)=\Pr(Y_{it}=1,Y_{i,t-1}=1\mid z_{st})\mathbf{\mathbbm{1}}\{z_{t}^{\prime}\beta_{0}$ $+\gamma_{0}\leq a\}$
$+\Pr(Y_{it}=1,Y_{i,t-1}=0\mid z_{st})\mathbf{\mathbbm{1}}\{z_{t}^{\prime}\beta_{0}\leq a\}.$
Now we discuss different scenarios for the value of $(z_{t}^{\prime}\beta_{0},z_{s}^{\prime}\beta_{0},\gamma_{0})$, leading to various identifying restrictions in Khan, Ponomareva, and Tamer (2023) described in Lemma 3.
Case 1: consider that $z_{t}^{\prime}\beta_{0}+\max\{0,\gamma_{0}\}\leq z_{s}^{\prime}\beta_{0}+\min\{0,\gamma_{0}\}$. Let $a=z_{t}^{\prime}\beta_{0}+\max\{0,\gamma_{0}\}$, the bounds can be simplified as
$U_{s}(a)$ $=1-\Pr(Y_{is}=0,Y_{i,s-1}=1\mid z_{st})-\Pr(Y_{is}=0,Y_{i,s-1}=0\mid z_{st})=\Pr(Y_{is}=1\mid z_{st}),$
$L_{t}(a)$ $=\Pr(Y_{it}=1,Y_{i,t-1}=1\mid z_{st})+\Pr(Y_{it}=1,Y_{i,t-1}=0\mid z_{st})=\Pr(Y_{it}=1\mid z_{st}).$
By exploiting the condition $\max\{0,\gamma_{0}\}-\min\{0,\gamma_{0}\}=|\gamma_{0}|$, the identifying condition is given as
$(z_{t}^{\prime}\beta_{0}-z_{s}^{\prime}\beta_{0})+|\gamma_{0}|\leq 0\Longrightarrow U_{s}(a)=\Pr(Y_{is}=1\mid z_{st})\geq L_{t}(a)=\Pr(Y_{it}=1\mid z_{st}).$
The contraposition of the above condition yields condition (i) in Lemma 3:
$\Pr(Y_{it}=1\mid z_{st})>\Pr(Y_{is}=1\mid z_{st})\Longrightarrow(z_{t}^{\prime}\beta_{0}-z_{s}^{\prime}\beta_{0})+|\gamma_{0}|>0.$
Case 2: consider that $z_{t}^{\prime}\beta_{0}+\max\{0,\gamma_{0}\}\leq z_{s}^{\prime}\beta_{0}+\gamma_{0}$. Let $a=z_{s}^{\prime}\beta_{0}+\gamma_{0}$, the bounds are simplified as
$U_{s}(a)\leq 1-\Pr(Y_{is}=0,Y_{i,s-1}=1\mid z_{st}),\quad L_{t}(a)=\Pr(Y_{it}=1\mid z_{st}).$
Given that $\max\{0,\gamma_{0}\}-\gamma_{0}=-\min\{0,\gamma_{0}\}$, we have the following identifying condition:
$z_{t}^{\prime}\beta_{0}-z_{s}^{\prime}\beta_{0}-\min\{0,\gamma_{0}\}\leq 0\Longrightarrow 1-\Pr(Y_{is}=0,Y_{i,s-1}=1\mid z_{st})\geq\Pr(Y_{it}=1\mid z_{st}).$
The contraposition of this condition leads to condition (ii) in Lemma 3:
$\Pr(Y_{it}=1\mid z_{st})>1-\Pr(Y_{is}=0,Y_{i,s-1}=1\mid z_{st})\Longrightarrow z_{t}^{\prime}\beta_{0}-z_{s}^{\prime}\beta_{0}-\min\{0,\gamma_{0}\}>0.$
The rest of conditions in Lemma 3 can be derived in a similar way so we omit the analysis here.
A.5 Proof of Proposition 2
@@
The identifying restriction (9) in Proposition 4 is given as follows:
$U_{t,order}(a\mid z_{st})$ $:=1-\sum_{j=1}^{J}\Pr(Y_{is}=y_{j},b_{j}-z_{s}^{\prime}\beta-X_{is}^{\prime}\gamma\geq a\mid z_{st})$ (18)
$\geq\sum_{j=1}^{J}\Pr(Y_{it}=y_{j},b_{j+1}-z_{t}^{\prime}\beta-X_{it}^{\prime}\gamma\leq a\mid z_{st}):=L_{t,order}(a\mid z_{st}).$
The proof idea of Corollary 2 is the same as Appendix 1 for Corollary 1. Let $q_{k}\in\mathcal{Q}_{2}$ denote a point in $\mathcal{Q}_{2}$, and we can rank the $2K(J-1)$ points as $q_{1}\leq q_{2}...\leq q_{2K(J-1)}$.
When $a>q_{2K(J-1)}$, then $\Pr(Y_{is}=y_{j},b_{j}-z_{s}^{\prime}\beta-X_{is}^{\prime}\gamma\geq a\mid z_{st})=0$ for any $j$ such that the upper bound is one $U_{t,order}(a\mid z_{st})=1$. Therefore, condition (18) holds for any $\theta$ and it is not informative for $\theta$.
When $a<q_{1}$, then $\Pr(Y_{it}=y_{j},b_{j+1}-z_{t}^{\prime}\beta-X_{it}^{\prime}\gamma\leq a\mid z_{st})=0$ for any $2\leq j\leq K$ so that the lower bound is zero $L_{t,order}(a\mid z_{st})=0$ and it is not informative for $\theta$.
When $a\in(q_{k},q_{k+1})$, we show that the identifying condition (18) at $q_{k+1}$ contains more information than the one at $a$. The lower bound $L_{t,order}(a\mid z_{st})$ is increasing in $a$, therefore the lower bound evaluated at $q_{k+1}$ is larger than the one evaluated at $a$. The upper bound is the same when evaluated at $q_{k+1}$ and at $a$, as there does not exist any point $b_{j}-z_{s}^{\prime}\beta-a_{k}^{\prime}\gamma$ that lies in $[a,q_{k+1})$ from the definition of $q_{k}$. Therefore, if the parameter $\theta$ satisfies condition (18) evaluated at $q_{k+1}$, it must satisfy condition (18) evaluated at $a$. The identifying condition at $a$ does not provide additional information for $\theta$.
A.6 Proof of Lemma 1

### v3

@@ A.4 Proof of Propositions 3 and 4
$1-\mathbb{P}(Y_{i1}=0,z_{1}^{\prime}\beta_{0}+X_{i1}^{\prime}\gamma_{0}\geq c\mid z)\geq\mathbb{P}(Y_{i2}=1,z_{2}^{\prime}\beta_{0}+X_{i2}^{\prime}\gamma_{0}\leq c\mid z),$
for $c\in\{z_{t}^{\prime}\beta_{0}+x_{k}^{\prime}\gamma_{0},\ t=1,2,k=1,...,K\}$, and another identifying condition switches the order of period 1 and 2.
Let $c=z_{1}^{\prime}\beta_{0}+x_{k}^{\prime}\gamma_{0}$, [^11] The value of $c=z_{2}^{\prime}\beta_{0}+x_{k}^{\prime}\gamma_{0}$ leads to the same identifying condition., then the above upper bound can be further bounded as
$1-\mathbb{P}(Y_{i1}=0,z_{1}^{\prime}\beta_{0}+X_{i1}^{\prime}\gamma_{0}\geq z_{1}^{\prime}\beta_{0}+x_{k}^{\prime}\gamma_{0}\mid z)\leq 1-\mathbb{P}(Y_{i1}=0,X_{i1}=x_{k}\mid z).$
When $z_{1}^{\prime}\beta_{0}-z_{2}^{\prime}\beta_{0}\geq 0$ which implies $z_{1}^{\prime}\beta_{0}+x_{k}^{\prime}\gamma_{0}\geq z_{2}^{\prime}\beta_{0}+x_{k}^{\prime}\gamma_{0}$, then the lower bound can be bounded below as
$\mathbb{P}(Y_{i2}=1,z_{2}^{\prime}\beta_{0}+X_{i2}^{\prime}\gamma_{0}\leq z_{1}^{\prime}\beta_{0}+x_{k}^{\prime}\gamma_{0}\mid z)\leq\mathbb{P}(Y_{i2}=1,X_{i2}=x_{k}\mid z).$
@@
If $\text{If }z_{1}^{\prime}\beta_{0}-z_{2}^{\prime}\beta_{0}\geq 0\Longrightarrow 1-\mathbb{P}(Y_{i1}=0,X_{i1}=x_{k}\mid z)\geq\mathbb{P}(Y_{i2}=1,X_{i2}=x_{k}\mid z).$
@@
$01-\mathbb{P}(Y_{i1}=0,X_{i1}=x_{k}\mid z)<\mathbb{P}(Y_{i2}=1,X_{i2}=x_{k}\mid z)\Longrightarrow\Delta z^{\prime}\beta_{0}>$ 0 .
@@
$01-\mathbb{P}(Y_{i1}=1,X_{i1}=x_{k}\mid z)<\mathbb{P}(Y_{i2}=0,X_{i2}=x_{k}\mid z)\Longrightarrow\Delta z^{\prime}\beta_{0}<$ 0 .
@@
$(z_{1},z_{2})\in\mathcal{Z}_{3}^{j}\ \Longrightarrow\ (x_{1}^{j}-x_{2}^{j})\gamma_{0}^{j}<\Delta z^{\prime}\beta_{0},$
$(z_{1},z_{2})\in\mathcal{Z}_{4}^{j}\ \Longrightarrow\ (x_{1}^{j}-x_{2}^{j})\gamma_{0}^{j}>\Delta z^{\prime}\beta_{0}.$
As previously shown, when $(z_{1},z_{2})\in\mathcal{Z}_{2}$, it implies that $\Delta z^{\prime}\beta_{0}<0$. Therefore, when $(z_{1},z_{2})\in\mathcal{Z}_{2}\cap\mathcal{Z}_{3}^{j}$, we have $(x_{1}^{j}-x_{2}^{j})\gamma_{0}^{j}<\Delta z^{\prime}\beta_{0}<0$ and the sign of $\gamma_{0}^{j}$ is identified given $x_{1}^{j}\neq x_{2}^{j}$. Similarly, when $(z_{1},z_{2})\in\mathcal{Z}_{1}\cap\mathcal{Z}_{4}^{j}$, the sign of $\gamma_{0}^{j}$ is also identified given $(x_{1}^{j}-x_{2}^{j})\gamma_{0}^{j}>\Delta z^{\prime}\beta_{0}>0$. Proposition 4 requires that for any $j\leq d_{x}$, either $\mathcal{Z}_{2}\cap\mathcal{Z}_{3}^{j}\neq\emptyset$ or $\mathcal{Z}_{1}\cap\mathcal{Z}_{4}^{j}\neq\emptyset$ so that the sign of $\gamma_{0}^{j}$ is identified for any $j$. ∎
A.5 Identification of Counterfactual Parameters
In the main text, we have focused on the (partial) identification of the index parameters $\theta_{0}$. Here we show how our identification results can also be leveraged to (partially) identify counterfactual parameters.
Write $W_{i}:=\left(Z_{i},X_{i}\right)$ in short, and correspondingly $w:=\left(z,x\right)$, and $w_{t}^{{}^{\prime}}\theta=z_{t}^{{}^{\prime}}\beta+x_{t}^{{}^{\prime}}\gamma$. Consider a general counterfactual change in the observable covariates $W_{i}$ from $w$ to $\tilde{w}$, and the consequent counterfactual period-$t$ conditional choice probability of the form
$\tilde{p}_{t}\left(\tilde{w}\right):=\mathbb{P}\left(\left.v_{it}\leq\tilde{w}^{{}^{\prime}}\theta_{0}\right|W_{i}=w\right).$ (50)
Importantly, in the definition above, the utility index is changed from $w^{{}^{\prime}}_{t}\theta_{0}$ to the counterfactual $\tilde{w}^{{}^{\prime}}_{t}\theta_{0}$, while the conditional distribution of the latent $v_{it}$ is held unchanged at $\left.v_{it}\right|W_{i}=w$. Hence, $\tilde{p}_{t}\left(\tilde{w}\right)$ can be interpreted as a counterfactual CCP induced by an exogenous policy intervention that only changes the characteristics from $w$ to $\tilde{w}$, but leaves all other unobserved individual heterogeneity reflected in the distribution of $v_{it}$ unchanged. In particular, note that the (partial) derivative of $\tilde{p}_{t}\left(w\right)$ can be interpreted as average marginal effects.
Proposition 10 (Bounds on Counterfactual CCP).
Under model 5 and Assumption 1,
$\inf_{\theta\in\Theta_{I}}\mathbb{P}\left(\left.Y_{it}=1,\ w_{t}^{{}^{\prime}}\theta\leq\tilde{w}_{t}^{{}^{\prime}}\theta\right|W_{i}=w\right)\leq\tilde{p}_{t}\left(\tilde{w}\right)\leq 1-\inf_{\theta\in\Theta_{I}}\mathbb{P}\left(\left.Y_{it}=0,\ w_{t}^{{}^{\prime}}\theta\geq\tilde{w}_{t}^{{}^{\prime}}\theta\right|W_{i}=w\right).$ (51)
@@ A.5 Identification of Counterfactual Parameters
By (52), we have
$\mathbb{P}\left(\left.Y_{it}=1,\ w_{t}^{{}^{\prime}}\theta_{0}\leq c\right|W_{i}=w\right)\leq F_{t}\left(\left.c\right|w\right)\leq 1-\mathbb{P}\left(\left.Y_{it}=0,\ w_{t}^{{}^{\prime}}\theta_{0}\geq c\right|W_{i}=w\right).$
Since $\tilde{p}_{t}\left(\tilde{w}\right)=F_{t}\left(\left.\tilde{w}_{t}^{{}^{\prime}}\theta_{0}\right|w\right)$, we have
$\mathbb{P}\left(\left.Y_{it}=1,\ w_{t}^{{}^{\prime}}\theta_{0}\leq\tilde{w}_{t}^{{}^{\prime}}\theta_{0}\right|W_{i}=w\right)\leq\tilde{p}_{t}\left(\tilde{w}\right)\leq 1-\mathbb{P}\left(\left.Y_{it}=0,\ w_{t}^{{}^{\prime}}\theta_{0}\geq\tilde{w}_{t}^{{}^{\prime}}\theta_{0}\right|W_{i}=w\right),$
and hence
$\inf_{\theta\in\Theta_{I}}\mathbb{P}\left(\left.Y_{it}=1,\ w_{t}^{{}^{\prime}}\theta\leq\tilde{w}_{t}^{{}^{\prime}}\theta\right|W_{i}=w\right)\leq\tilde{p}_{t}\left(\tilde{w}\right)\leq 1-\inf_{\theta\in\Theta_{I}}\mathbb{P}\left(\left.Y_{it}=0,\ w_{t}^{{}^{\prime}}\theta\geq\tilde{w}_{t}^{{}^{\prime}}\theta\right|W_{i}=w\right).$
The lower and upper bounds in Proposition 10 above are identified since the involved conditional probabilities are all about observed data $\left(Y_{i},W_{i}\right)$ for each $\theta\in\Theta_{I}$, while the set $\Theta_{I}$ is identified by Proposition 1. Hence, Proposition 10 establishes the partial identification of the counterfactual CCP $\tilde{p}_{t}\left(\tilde{w}\right)$.
Note that the (partial) identification of counterfactual CCP $\tilde{p}_{t}\left(\tilde{w}\right)$ relies on the identification of the index parameter $\theta_{0}$ as well as the identification of the latent conditional distribution $\left.v_{it}\right|W_{i}=w$, which also involves the endogenous covariates $X_{i}$. It turns out that, our key identification strategy in Section 2.2 also provides a straightforward way to derive bounds on $F_{t}\left(\left.c\right|w\right)$, the CDF of $\left.v_{it}\right|W_{i}=w$ at any point $c$, by taking conditional expectations of (7) and (8) given $W_{i}=w$ (instead of $Z_{i}=z$ as in Section 2.2):
$\mathbb{P}\left(\left.Y_{it}=1,\ w_{t}^{{}^{\prime}}\theta_{0}\leq c\right|W_{i}=w\right)\leq F_{t}\left(\left.c\right|w\right)\leq 1-\mathbb{P}\left(\left.Y_{it}=0,\ w_{t}^{{}^{\prime}}\theta_{0}\geq c\right|W_{i}=w\right),$ (52)
which can then be combined with Proposition 1 to derive the bounds in Proposition 10.
A.6 Proof of Proposition 8
@@
Let $v_{ijt}:=\alpha_{ij}+\epsilon_{ijt}$, for any set $K\subset\mathcal{J}$, the probability of selecting a choice $j\in K$ conditional on $W_{i}=w$ is given as:
$\mathbb{P}(Y_{it}^{K}\mid w)=\mathbb{P}(Y_{it}\in K\mid w)=\mathbb{P}\left(\exists j\in K\text{ s.t. }w_{ijt}^{\prime}\theta_{0}+v_{ijt}\geq w_{ikt}^{\prime}\theta_{0}+v_{ikt}\ \forall k\in K^{c}\mid w\right).$
The above observed probability restricts the conditional distribution of $v_{ikt}-v_{ijt}\mid w$ and can be exploited to bound this distribution.
We define $Q_{t}(c_{jk}\mid w)$ as follows: for $c_{jk}\in\mathcal{R}$,
$Q_{t}(c_{jk}\mid w):=\mathbb{P}\left(\exists j\in K\text{ s.t. }v_{ikt}-v_{ijt}\leq c_{jk}\ \forall k\in\mathcal{J}\setminus K\mid w\right).$
Then, we can derive lower and upper bounds for the above probability using variations in observed choice probabilities. When $c_{jk}$ satisfies $c_{jk}\geq(w_{ijt}-w_{ikt})^{\prime}\theta_{0}$ for any $j\in K$ and $k\in\mathcal{J}\setminus K$, then $Q_{t}(c_{jk}\mid w)$ can be bounded below as
$Q_{t}(c_{jk}\mid w)$ $\geq\mathbb{P}\left(\exists j\in K\text{ s.t. }v_{ikt}-v_{ijt}\leq(w_{ijt}-w_{ikt})^{\prime}\theta_{0}\ \forall k\in\mathcal{J}\setminus K\mid w\right)$
$=\mathbb{P}(Y_{it}\in K\mid w).$
Therefore, the lower bound for $Q_{t}(c_{jk}\mid w)$ is established as
$Q_{t}(c_{jk}\mid w)\geq\mathbb{P}(Y_{it}\in K,\ c_{jk}\geq(w_{ijt}-w_{ikt})^{\prime}\theta_{0}\ \forall j\in K,k\in k\in\mathcal{J}\setminus K\mid w).$
The above inequality holds since either $c_{jk}\geq(w_{ijt}-w_{ikt})^{\prime}\theta_{0}$ or the lower bound is zero.
By taking expectation of $X_{i}$ given $z$, we can bound the conditional distribution $Q_{t}(c_{jk}\mid z)$ as
$Q_{t}(c_{jk}\mid z)$ $\geq\mathbb{P}\left(Y_{it}\in K,c_{jk}\geq(z_{ijt}-z_{ikt})^{\prime}\beta_{0}+(X_{ijt}-X_{ikt})^{\prime}\gamma_{0}\ \forall j\in K,k\in\mathcal{J}\setminus K\mid z\right)$
$=\mathbb{P}\left(Y_{it}^{K}=1,c_{jk}\geq(z_{ijt}-z_{ikt})^{\prime}\beta_{0}+(X_{ijt}-X_{ikt})^{\prime}\gamma_{0}\ \forall j\in K,k\in\mathcal{J}\setminus K\mid z\right).$
Similarly, the conditional probability $Q_{t}(c_{jk}\mid w)$ can be bounded above as
$Q_{t}(c_{jk}\mid w)\leq\mathbb{P}(Y_{it}^{K}=1\mid w)\mathbf{\mathbbm{1}}\{c_{jk}\leq(w_{ijt}-w_{ikt})^{\prime}\theta_{0}\ \forall j\in K,\mathcal{J}\setminus K\}+\\ 1-\mathbf{\mathbbm{1}}\{c_{jk}\leq(w_{ijt}-w_{ikt})^{\prime}\theta_{0}\ \forall j\in K,k\in\mathcal{J}\setminus K\}.$
The above inequality holds since either $c_{jk}\leq(w_{ijt}-w_{ikt})^{\prime}\theta_{0}$ or the upper bound is one with $c_{jk}>(w_{ijt}-w_{ikt})^{\prime}\theta_{0}$. After taking expectation of $X_{i}$ given $z$, the upper bound for $Q_{t}(c_{jk}\mid z)$ is obtained as
$Q_{t}(c_{jk}\mid z)\leq\mathbb{P}\left(Y_{it}^{K}=1,c_{jk}\leq(z_{ijt}-z_{ikt})^{\prime}\beta_{0}+(X_{ijt}-X_{ikt})^{\prime}\gamma_{0}\ \forall j\in K,k\in K^{c}\mid z\right)\\ +1-\mathbb{P}\left(c_{jk}\leq(z_{ijt}-z_{ikt})^{\prime}\beta_{0}+(X_{ijt}-X_{ikt})^{\prime}\gamma_{0}\ \forall j\in K,k\in\mathcal{J}\setminus K\mid z\right).$
Rearranging the above formula yields
$Q_{t}(c_{jk}\mid z)\leq 1-\mathbb{P}\left(Y_{it}^{K}=0,c_{jk}\leq(z_{ijt}-z_{ikt})^{\prime}\beta_{0}+(X_{ijt}-X_{ikt})^{\prime}\gamma_{0}\ \forall j\in K,k\in\mathcal{J}\setminus K\mid z\right).$
Under Assumption 1, the conditional probability $Q_{t}(c_{jk}\mid z)$ is the same for any $t$. Therefore, the smallest upper bound of $Q_{t}(c_{jk}\mid z)$ should be larger than the largest lower bound over all periods, yielding the identifying condition (23) as follows:
$1-\max_{s=1,...,T}\mathbb{P}(Y_{is}^{K}=0,(z_{js}-z_{ks})^{\prime}\beta_{0}+(X_{ijs}-X_{iks})^{\prime}\gamma_{0}\geq c_{jk}\ \forall j\in K,k\in\mathcal{J}\setminus K\mid z)\\ \geq\max_{t=1,...,T}\mathbb{P}(Y_{it}^{K}=1,(z_{jt}-z_{kt})^{\prime}\beta_{0}+(X_{ijt}-X_{ikt})^{\prime}\gamma_{0}\leq c_{jk}\ \forall j\in K,k\in\mathcal{J}\setminus K\mid z).$
A.7 Proof of Proposition 9